# Model Merging in LLMs, MLLMs, and Beyond: Methods, Theories, Applications and Opportunities

### v4

@@ 3.1.5 Combine the Capabilities of Expert LLMs
3.2 Model Merging in Multimodal Large Language Models (MLLMs)
@@
An interesting question is whether we can merge multiple modality-specific models to obtain a single, effective, and parameter-efficient modality-agnostic model. We aim for the merged unified model to encode inputs from different modalities, learn cross-modal interactions, and maintain performance comparable to that of well-trained independent modality-specific models. Compared to traditional multimodal learning, model merging techniques offer new opportunities. This model-merging approach offers several benefits: (1) it eliminates the costly and labor-intensive process of collecting labeled paired multimodal training examples, which is required for jointly training multimodal models; (2) it enhances the adaptability of multimodal models, allowing for the seamless integration of new modalities; and (3) it fully leverages knowledge collaboration across multiple modalities, thereby benefiting from cross-modal knowledge transfer.
3.2.1 Model Merging for Multimodal Fusion
@@
Recently, many studies have focused on merging models from different modalities into a single model, thereby enhancing the diversity of knowledge across modalities. For instance, JAM [4] proposes to merge two specialized (one for text-to-image and one text-only) autoregressive, decoder-only, large transformer models to seamlessly generate multimodal outputs. Similarly, DAMC [18] introduces a method for fusing multimodal LLMs across image, audio, video, and point cloud modalities, further reducing cross-modal interference through parameter decoupling and adjusting modality fusion coefficients.
@@
To evaluate the impact of various factors on model merging, VL-Merging [174] performs a comprehensive empirical analysis of multimodal model merging. The overall framework consists of three steps: independent modality fine-tuning, multimodal merging, and downstream task fine-tuning. Through experiments involving different initializations, merging methods, and architectures in multimodal model merging, the authors propose the following guidelines: (1) Models across multiple modalities should be based on the same pretraining starting point to ensure they are in the same basin [5] and share more information. (2) Simple model averaging achieves better performance, and if more computation and storage resources are available, more fine-grained merges can be conducted. (3) Merging the entire model rather than just a subset of layers generally yields more satisfactory results, as fine-tuning only a subset of layers may restrict the capabilities of single-modality models. Unlike the above model merging approaches that are developed based on specific architectures, UnIVAL [165] is the first to design a unified architecture for four modalities: image, video, audio, and language. It transforms the tasks across all modalities into a “sequence-to-sequence” format, with the training objectives of all modalities converted into a “next token prediction” format. This allows for a uniform feature extraction and classifier to be applied across all modalities. Additionally, UnIVAL provides favorable architectural conditions for model merging and demonstrates that linear interpolation of models fine-tuned across multiple modalities in weight space results in a general single model that performs well on both seen and unseen tasks.
3.2.2 Model Merging for Cross-Modal Knowledge Transfer
@@
Some works attempt to transfer knowledge from one modality to another through a model merging approach. For instance, MAM [173] investigates whether the attention layers of Transformers [191] generalize across different modalities. Specifically, it examines if the knowledge acquired by Transformer models trained on high-resource modalities (e.g., data-rich images and text) can be transferred to Transformer models trained on low-resource modalities (e.g., data-sparse speech and audio). This paper demonstrates attention merging for models across various tasks, modalities, and initializations. The final results show that MAM achieves an 18.42% reduction in classification error on the audio classification task (using the ESC-50 dataset [142]) compared to the standard fine-tuning paradigm.
3.3 Model Merging in Visual Generative Models
@@
The goal of visual generative models, such as generative adversarial networks (GANs), variational autoencoders (VAEs), normalizing flows (Flows), and denoising diffusion probabilistic models (Diffusions), is to approximate the underlying data distribution behind a given dataset so as to generate more new samples with the same distribution. However, image generative models still face the following challenges: the inability to flexibly generate samples with multiple style combinations, the high cost of generative model training, and the inability to generate all the details specified in the instructions. This dilemma has led to an interest in expert models, which train a set of experts with specific abilities on different data shards or distributions, allowing for the flexible addition or removal of certain styles of experts at inference time. Considering the difficulty in deploying and the cost of resources for ensemble learning, model merging offers a new perspective on combining skill-specific experts of different styles without additional memory and inference costs.
3.3.1 Style Mixing in Generative Models
Existing generative models typically generate distributions based only on the training data. However, in real deployments, different users or artists often want to generate artwork with different combinations of styles. Collecting additional data for these mixed distributions is expensive, and fine-tuning the model can result in the forgetting of other capabilities. Model merging offers the potential to flexibly combine multiple styles.
@@ 3.3.1 Style Mixing in Generative Models
Earl GAN Cocktail [10] attempted to merge several pre-trained GAN models. Recently, diffusion-based image generative models [62, 156, 157] have gained more attention than GAN-based models due to their superior generative capabilities. Consequently, most research focuses on fusing different diffusion models. Specifically, Diffusion Soup [14] demonstrates the ability to linearly merge diffusion models fine-tuned on data shards of different styles (e.g., data provided by different domains/categories or different users), resulting in hybrid style zero-shot generation. In addition, Diffusion Soup empirically verifies that model merging has an anti-memorization effect, meaning the generated images are less likely to replicate training data, which is beneficial for generating diverse images. Unlike Diffusion Soup, which directly merges model parameters, MaxFusion [132] is inspired by ZipIt [170] and proposes merging intermediate features of multiple diffusion models based on the same input noise to generate images that satisfy multiple conditions. However, merging multiple diffusion models based on full parameter fine-tuning can be costly when the number of tasks is large. To address this issue, ZipLoRA [163] and MoLE [212] aim to seamlessly merge parameter-efficient LoRA modules. For example, ZipLoRA proposes merging independently trained content/subject (e.g., a specific object or person) LoRAs with artistic style (e.g., drawing or painting, etc.) LoRAs, allowing the diffusion model to generate any user-provided combination of subjects and styles [158]. This approach enables users and artists to easily combine publicly available subjects and styles LoRAs of their choice.
3.3.2 Reducing Training Cost of Generative Models
[…]
3.3.3 Enhancing the Faithfulness of Generative Models
Some studies on Text-to-Image (T2I) show that although the existing T2I generation model can generate high-quality images according to text prompts, the images often fail to fully capture and reflect the semantic details in the text, such as generating multiple subjects or correctly depicting the spatial relationships between objects [101]. To enhance the fidelity of the T2I generation models, SELMA [101] designed a novel four-stage paradigm. In the first and second stages, a series of input texts (corresponding to different skills) are collected through diversified prompts from existing LLMs, and the corresponding image data are generated using the T2I model. The third stage involves fine-tuning skill-specific experts (i.e., LoRA) separately on images of different skills. In the fourth stage, expert models with different skills are merged to obtain the final model during inference. Compared to the paradigm of joint learning of multiple skills, this method of merging expert skills after independent learning may help alleviate knowledge/skill conflicts while also being more efficient.
4 Application of Model Merging in Different Machine Learning Subfields

### v5

@@ 4.1. Model Merging in Large Language Models (LLMs)
Figure 7. Example application scenarios of model merging in large language models.
4.1.1. Human Preference Alignment for LLMs
Humans often hold diverse opinions about aesthetics, politics, or fairness. When LLMs serve humans, different people have different expectations of the model, e.g., some expect LLMs to generate harmless responses, while others seek engaging and enjoyable interactions (Rame et al., 2023). Consequently, the development of practical LLMs is generally divided into three stages, to generate responses that are more helpful, accurate, and safer (Lu et al., 2024b): Pre-training on a large amount of unsupervised data, supervised fine-tuning (SFT) on a small dataset with high-quality annotation, and interaction with humans to further optimize LLM alignment (e.g., direct preference optimization (DPO) (Rafailov et al., 2023) or reinforcement learning from human feedback (RLHF) (Ziegler et al., 2019)) with human preferences, rewards or values.
Some works propose to achieve better, safer, or faster alignment of human preferences by model merging (Zheng et al., 2024; Wang et al., 2024b; Bhardwaj et al., 2024; Fu et al., 2024; Pentyala et al., 2024). For example, ExPO (Zheng et al., 2024) adds a task vector, constructed by a moderate model aligned using DPO or RLHF on a small amount of human preference data, to an unaligned SFT model. A more powerful aligned model can be directly obtained by setting a suitable merging coefficient. On the AlpacaEval 2.0 benchmark (Li et al., 2023d), fusing a model aligned on the 10%/20% preference data with an SFT model results in performance comparable to that of a model aligned on the full preference data. DogeRM (Lin et al., 2024) proposed merging the reward model with LLMs fine-tuned on different downstream domains to create domain-private reward models directly. Additionally, Lu et al. (2024b) propose an Online Merging Optimizer, that interpolates the gradient with the SFT model at each step of RLHF. This approach encourages RLHF to optimize toward reward maximization while preventing LLMs from forgetting general knowledge due to RLHF. Beyond preference alignment, several studies have examined the impact of model merging for secure alignment of LLMs (Bhardwaj et al., 2024; Yi et al., 2024; Hammoud et al., 2024; Hazra et al., 2024). For example, Hammoud et al. (2024) find that merging two security-aligned models could compromise security. Thus, they proposed explicitly including secure alignment as an optimization objective when constructing synthetic data for model merging.
In practice, users often have various combinations of preferences rather than a single preference. Training a model separately for each combination of preferences is unrealistic due to the infinite combinations and the high training costs. Therefore, some studies suggest combining models with different reward alignments to create a series of integrated aligned LLMs. For example, Rame et al. (2023) and Jang et al. (2023a) propose Reward Soups and Personalized Soups, respectively, as efficient and flexible solutions for diverse rewards. Specifically, Rewarded Soups first trains an expert model for each reward and then linearly interpolates the weights of the experts to approximate the set of Pareto optimal solutions for various reward combinations. This approach is cost-effective, as it only requires training separate models for each reward to combine any variety of rewards.
4.1.2. Detoxification of LLMs
LLMs have been widely noted for issues related to untruthfulness and toxicity in various applications (Hu et al., 2024), such as insults, threats, and profanity in responses to certain questions. To address the potential security risks in the application of LLMs, flexible techniques are needed to reduce the generation of toxic text, essentially detoxifying LLMs. A straightforward solution is to collect additional non-toxic data to fine-tune LLMs (Keskar et al., 2019); however, this approach requires significant computing resources and may interfere with the general capabilities of LLMs. Alternatively, directly reducing the probability of potentially toxic words during the decoding stage requires additional guidance information (Krause et al., 2021). Recent studies have shown that reducing the toxic data generation of LLMs through model merging is a simple and effective scheme (Ilharco et al., 2023; Zhang et al., 2023a; Hu et al., 2024; Dige et al., 2024).
Task Arithmetic (Ilharco et al., 2023) negates the task vectors of GPT-2 model (Radford et al., 2019) fine-tuned on toxic data (Civil Comments (Borkan et al., 2019)) and shows that this operation effectively reduces the proportion of data classified as ”toxic”, with little change in the fluency of the language on the control task (WikiText-103). Additionally, some parameter-efficient models steer the toxic behavior of LLMs by manipulating a small number of parameters. PEM (Zhang et al., 2023a) negates LoRA (Hu et al., 2022) (and (IA)3 (Liu et al., 2022b)) modules trained on poisoning data to maintain language proficiency while reducing the toxicity of language model output. Ethos (Gao et al., 2024) and Ext-Sub (Hu et al., 2024) point out that while the task vector on toxic data is factually wrong, it also contains correct information about language modeling and logical narrative skills. Therefore, Ext-Sub decomposes the toxic task vector into two orthogonal subspaces that represent general capability and destructive capability, respectively. Toxic knowledge is then eliminated by removing only the component representing the destructive ability from the LLM.
4.1.3. Knowledge Unlearning of LLMs
LLMs may inadvertently learn copyrighted material, raising significant legal and ethical concerns (Abad et al., 2024), and broader questions about responsible AI use (Dou et al., 2024). In this context, the California Consumer Privacy Act (Pardau, 2018) and the General Data Protection Regulations of the European Union (Hoofnagle et al., 2019) stipulate the right to data forgetting. The foundational model’s knowledge must be adapted to comply with these regulations. However, the cost of excluding copyrighted data for retraining from scratch is prohibitive. For instance, training a Llama-2-70B from scratch requires 1,720,320 GPU hours (Touvron et al., 2023b). Traditional methods often use gradient ascent (GA) to achieve forgetting by fine-tuning the model using the GA algorithm on the specific data to be forgotten (Thudi et al., 2022; Yao et al., 2023). Unfortunately, this approach typically catastrophically destroys other parts of the model’s knowledge. That is, forgetting specific knowledge also erases other knowledge that should be retained. Recently, many studies based on model merging techniques have demonstrated the potential to forget LLM-specific knowledge without harming other knowledge (Ilharco et al., 2023; Dou et al., 2024; Hu et al., 2024).
Unlike the GA-based approach, the model merging approach does not require additional data for other tasks to maintain old knowledge. To achieve forgetting, model merging typically incorporates a negatively fine-tuned model into the target model (i.e., the task-specific fine-tuned knowledge is subtracted from the target model). For example, Task Arithmetic (Ilharco et al., 2023) shows that negating task vectors degrades performance on specific tasks without substantial changes to the control tasks. Experiments demonstrate that model merging can forget the knowledge of the target task in a fine-tuned model without harming performance on control tasks. Similarly, Stable Sequential Unlearning (Dou et al., 2024) extends this forgetting to the setting of sequential unlearning on LLMs, where different copyrighted content must be unlearned at different time steps. Knowledge forgetting can also forget samples that represent bad behavior during pretraining. For instance, FuseToForget (Zaman et al., 2023) employs model merging as a debiasing tool to reduce privacy issues in language models. FLearning (Ni et al., 2023) first subtracts the parameters related to the data to be forgotten and then fine-tunes the parameters with new data to achieve accurate knowledge updates. SKU (Liu et al., 2024a) explores the forgetting of harmful data in LLM, which is a two-stage scheme. Initially, harmful data (e.g., harmful question-answer pairs) is used to fine-tune the parameters corresponding to the location of harmful knowledge in the LLM (i.e., the task vector), and then the task vector is negated from the LLM to mitigate undesirable behavior in the LLM effectively. Generally, incorporating the opposite (anti-expert) task vectors into the pre-trained model can effectively accomplish the task of machine unlearning.
4.1.4. Faster Training of LLMs
The training of LLMs requires numerous iterations on massive data, making the training process extremely expensive (Kaddour, 2022; Ram et al., 2024). For example, training LLAMA2-70B with 2T tokens required 1,720,320 GPU hours (Liu et al., 2024d). Methods to accelerate LLM training include mixed-precision training, continual retraining, and pipeline parallelism. An orthogonal approach is checkpoint merging in training trajectories, which offers a simple and effective means to either speed up LLM training or enhance training performance at the same cost. Some works incorporate checkpoints in a single training trajectory during LLM training to accelerate model training. For instance, LAWA (Kaddour, 2022) demonstrated that merging checkpoints during intermediate stages of model training speeds up the process. For example, training a RoBERTa-Base model on the WikiText-103 dataset saved 30 GPU hours. Sanyal et al. (2023) further showed that the combination of checkpoint averaging in pre-trained trajectories and a high learning rate contributes to faster convergence. Checkpoint Merging (Liu et al., 2024d) comprehensively evaluates the effectiveness of model merging at different stages of the Baichuan2 (Yang et al., 2023b) pre-training process. PMA (Li et al., 2025a) and WSM (Tian et al., 2025b) show that applying checkpoint averaging during LLM pre-training can serve as an effective alternative to learning rate decay schedules, while substantially reducing training costs.
4.1.5. Faster Inference of MoE-based LLMs
[…]
4.1.6. Faster Reasoning of LLMs
Reasoning-oriented LLMs (e.g., DeepSeek-R1 (Guo et al., 2025), QwQ-32B-Preview (Team, 2025b)) have achieved remarkable improvements on complex tasks by engaging in iterative, chain-of-thought reasoning. However, this enhanced reasoning capability comes at the cost of significantly lower efficiency compared with conventional LLMs, and in some cases, the model may overthink even very simple problems, consuming a large number of tokens. Prior work (Chen et al., 2024c) reports that QwQ-32B-Preview uses 901 tokens to answer a simple question such as “2+3=?”, whereas conventional models like Llama-3.3-70B and GPT-4o require only 7 tokens, and Gemini Pro only 5 tokens. Recently, a line of work has begun to reduce token consumption by merging “slow-thinking” reasoning models with “fast” conventional LLMs (Wu et al., 2025a; Team, 2025a; Wu et al., 2025b). For example, the Kimi team (Team, 2025a) finds that merging long-context and short-context models is highly effective in improving token efficiency. Wu et al. (2025b) conduct a detailed evaluation of different model merging methods and show that model merging provides an efficient mechanism for implementing long-to-short reasoning. Task-vector-based approaches, such as Task Arithmetic and TIES-Merging, can reduce inference length / token consumption by around 50% on DeepSeek-R1-7B and Qwen2.5-Math-7B models while maintaining comparable accuracy.
4.1.7. Combine the Capabilities of Expert LLMs
LLMs exhibit strong generalizability in general tasks, but often lack knowledge in specific vertical domains. Pretrained LLMs typically require fine-tuning within different corporations to become expert LLMs in various fields. Integrating the expertise of multiple specialists is particularly critical for solving more complex tasks (Tam et al., 2024b). Research on model merging techniques indicates that a composite LLM can be created by combining the parameters of different expert LLMs (Yu et al., 2024b; Akiba et al., 2025; Dekoninck et al., 2024; Zhou et al., 2024b; Wan et al., 2024a, b; Yu et al., 2024a; Sukhbaatar et al., 2024; Yu et al., 2024a; Li et al., 2024a; Cao et al., 2025). For example, Dekoninck et al. (2024) demonstrate the ability to flexibly control text generation by merging multiple LLMs with different styles and applying personalized weighting. Robust Weight Signatures (Cai et al., 2023) proposes a robustness “patching” framework via model merging to enhance the overall robustness of the model against various naturally corrupted versions of clean data. Kwai Keye-VL 1.5 (Yang et al., 2025b) leverages model merging to combine domain-specific expert models with a LongCoT cold-start model into a unified general model, leading to improved performance. Souper-Model (Maiti et al., 2025) demonstrates that model merging can improve both performance and stability across multiple capabilities, including multilingual understanding, tool calling, and mathematical reasoning, and achieves state-of-the-art results on the Berkeley Function Calling Leaderboard (Patil et al., 2025). In summary, model merging offers a straightforward and effective strategy for enhancing LLM’s capabilities.
4.2. Model Merging in Multimodal Large Language Models (MLLMs)
@@
An interesting question is whether we can merge multiple modality-specific models to obtain a single, effective, and parameter-efficient modality-agnostic model. As shown in Figure 8(left), we aim for the merged unified model to encode inputs from different modalities, learn cross-modal interactions, and maintain performance comparable to that of well-trained independent modality-specific models. Compared to traditional multimodal learning, model merging techniques offer new opportunities. This model-merging approach offers several benefits: (1) it eliminates the costly and labor-intensive process of collecting labeled paired multimodal training examples, which is required for jointly training multimodal models; (2) it enhances the adaptability of multimodal models, allowing for the seamless integration of new modalities; and (3) it fully leverages knowledge collaboration across multiple modalities, thereby benefiting from cross-modal knowledge transfer.
Figure 8. Example application scenarios of model merging in multimodal large language models (left) and visual generative models (right).
4.2.1. Model Merging for Multimodal Fusion
@@
Recently, many studies have focused on merging models from different modalities into a single model, thereby enhancing the diversity of knowledge across modalities. For instance, JAM (Aiello et al., 2024) proposes to merge two specialized (one for text-to-image and one text-only) autoregressive, decoder-only, large transformer models to seamlessly generate multimodal outputs. Similarly, DAMC (Chen et al., 2024a) introduces a method for fusing multimodal LLMs across image, audio, video, and point cloud modalities, further reducing cross-modal interference through parameter decoupling and adjusting modality fusion coefficients.
@@
To evaluate the impact of various factors on model merging, VL-Merging (Sung et al., 2023) performs a comprehensive empirical analysis of multimodal model merging. The overall framework consists of three steps: independent modality fine-tuning, multimodal merging, and downstream task fine-tuning. Through experiments involving different initializations, merging methods, and architectures in multimodal model merging, the authors propose the following guidelines: (1) Models across multiple modalities should be based on the same pretraining starting point to ensure they are in the same basin (Ainsworth et al., 2023) and share more information. (2) Simple model averaging achieves better performance, and if more computation and storage resources are available, more fine-grained merges can be conducted. (3) Merging the entire model rather than just a subset of layers generally yields more satisfactory results, as fine-tuning only a subset of layers may restrict the capabilities of single-modality models.
Unlike the above model merging approaches that are developed based on specific architectures, UnIVAL (Shukor et al., 2023) is the first to design a unified architecture for four modalities: image, video, audio, and language. It transforms the tasks across all modalities into a ”sequence-to-sequence” format, with the training objectives of all modalities converted into a ”next token prediction” format. This allows for a uniform feature extraction and classifier to be applied across all modalities. Additionally, UnIVAL provides favorable architectural conditions for model merging and demonstrates that linear interpolation of models fine-tuned across multiple modalities in weight space results in a general single model that performs well on both seen and unseen tasks.
4.2.2. Model Merging for Cross-Modal Knowledge Transfer
@@
Some works attempt to transfer knowledge from one modality to another through a model merging approach. For instance, MAM (Sundar et al., 2024) investigates whether the attention layers of Transformers (Vaswani et al., 2017) generalize across different modalities. Specifically, it examines if the knowledge acquired by Transformer models trained on high-resource modalities (e.g., data-rich images and text) can be transferred to Transformer models trained on low-resource modalities (e.g., data-sparse speech and audio). This paper demonstrates attention merging for models across various tasks, modalities, and initializations. The final results show that MAM achieves an 18.42% reduction in classification error on the audio classification task (using the ESC-50 dataset (Piczak, 2015)) compared to the standard fine-tuning paradigm.
4.3. Model Merging in Visual Generative Models
@@
The goal of visual generative models, such as generative adversarial networks (GANs) and denoising diffusion probabilistic models (Diffusions), is to approximate the underlying data distribution behind a given dataset so as to generate more new samples with the same distribution. However, image generative models still face the following challenges: the inability to flexibly generate samples with multiple style combinations and the high cost of generative model training. This dilemma has led to an interest in expert models, which train a set of experts with specific abilities on different data shards or distributions, allowing for the flexible addition or removal of certain styles of experts at inference time. Considering the difficulty in deploying and the cost of resources for ensemble learning, model merging offers a new perspective on combining skill-specific experts of different styles without additional memory and inference costs.
4.3.1. Style Mixing in Generative Models
@@
Existing generative models typically generate distributions based only on the training data. However, in real deployments, different users or artists often want to generate artwork with different combinations of styles. Collecting additional data for these mixed distributions is expensive, and fine-tuning the model can result in the forgetting of other capabilities. As shown in Figure 8(right), model merging offers the potential to flexibly combine multiple styles. Earl GAN Cocktail (Avrahami et al., 2022) attempted to merge several pre-trained GAN models. Recently, diffusion-based image generative models (Ho et al., 2020; Rombach et al., 2022) have gained more attention than GAN-based models due to their superior generative capabilities. Consequently, most research focuses on fusing different diffusion models. Specifically, Diffusion Soup (Biggs et al., 2024) demonstrates the ability to linearly merge diffusion models fine-tuned on data shards of different styles (e.g., data provided by different domains/categories or different users), resulting in hybrid style zero-shot generation. In addition, Diffusion Soup empirically verifies that model merging has an anti-memorization effect, meaning the generated images are less likely to replicate training data, which is beneficial for generating diverse images. Unlike Diffusion Soup, which directly merges model parameters, MaxFusion (Nair et al., 2024) is inspired by ZipIt (Stoica et al., 2024a) and proposes merging intermediate features of multiple diffusion models based on the same input noise to generate images that satisfy multiple conditions. However, merging multiple diffusion models based on full parameter fine-tuning can be costly when the number of tasks is large. To address this issue, ZipLoRA (Shah et al., 2023) and MoLE (Wu et al., 2024) aim to seamlessly merge parameter-efficient LoRA modules. For example, ZipLoRA proposes merging independently trained content/subject (e.g., a specific object or person) LoRAs with artistic style (e.g., drawing or painting, etc.) LoRAs, allowing the diffusion model to generate any user-provided combination of subjects and styles (Ryu, [n. d.]). This approach enables users and artists to easily combine publicly available subjects and styles LoRAs of their choice. Beyond image generation, Tian et al. (2025a) apply model merging to image-to-video generation, enabling synthesized video clips that not only exhibit natural motion but also preserve the original appearance of the input image.
4.3.2. Reducing Training Cost of Generative Models
In real-world scenarios, large-scale training data typically originates from different domains or is provided by various users. Given the need to add new data or remove outdated data, retraining a single model with updated data every time is often impractical (Biggs et al., 2024). For instance, training a CM model (Song et al., 2023) using 8 A100 GPUs (Liu et al., 2024e) takes about one week. This is because existing methods only apply the final convergence weights during generative model training and ignore the intermediate training trajectories (Thakur and Vashisth, 2024; Liu et al., 2024e). LCSC (Liu et al., 2024e) demonstrates that a simple combination of training trajectories in the middle of the diffusion model by the evolutionary algorithm can significantly reduce the training cost. Specifically, only a few iterations or a small batch size is required to train the diffusion model to achieve image quality comparable to that of a fully trained diffusion model. For example, on the CIFAR-10 dataset, LCSC improves the training process for consistency distillation and consistency training (Song et al., 2023) by factors of $\small 23\times$ and $\small 7\times$, respectively. The underlying reason is that each local checkpoint of the optimized trajectory has many high-quality basins (i.e., areas of better generation quality) nearby that cannot be reached by stochastic gradient descent due to substantial variance in gradient estimations. However, checkpoint interpolation provides an opportunity to reach these basins.
5. Application of Model Merging in Different Machine Learning Subfields